# OpenGaussian: Towards Point-Level 3D Gaussian-based Open Vocabulary Understanding

### v1

@@ Abstract
This paper introduces OpenGaussian, a method based on 3D Gaussian Splatting (3DGS) capable of 3D point-level open vocabulary understanding. Our primary motivation stems from observing that existing 3DGS-based open vocabulary methods mainly focus on 2D pixel-level parsing. These methods struggle with 3D point-level tasks due to weak feature expressiveness and inaccurate 2D-3D feature associations. To ensure robust feature presentation and 3D point-level understanding, we first employ SAM masks without cross-frame associations to train instance features with 3D consistency. These features exhibit both intra-object consistency and inter-object distinction. Then, we propose a two-stage codebook to discretize these features from coarse to fine levels. At the coarse level, we consider the positional information of 3D points to achieve location-based clustering, which is then refined at the fine level. Finally, we introduce an instance-level 3D-2D feature association method that links 3D points to 2D masks, which are further associated with 2D CLIP features. Extensive experiments, including open vocabulary-based 3D object selection, 3D point cloud understanding, click-based 3D object selection, and ablation studies, demonstrate the effectiveness of our proposed method. The source code will be released at our project page https://3d-aigc.github.io/OpenGaussian.
@@ 3 Method
3DGS [1] utilizes an explicit scene representation through 3D Gaussian points. Each Gaussian point encompasses various attributes such as position $\boldsymbol{\mu}$, rotation $\boldsymbol{R}$, scale $\boldsymbol{S}$, opacity $\sigma$, and spherical harmonics coefficients for representing direction-aware color $\boldsymbol{c}$. For a detailed description of the splatting process, please refer to Appendix C, where a set of 3D Gaussian points is projected onto 2D screen space and blended to generate pixels. Inspired from prior studies [14, 15, 16], we augment each 3D Gaussian point with a low-dimensional feature $\boldsymbol{f}\in\mathbb{R}^{6}$ to represent its instance attributes. However, our approach differs in two crucial aspects: 1) We do not require additional dimensional reduction, quantization, or distillation for pre-trained features (such as CLIP [22], SAM [23], DINO [42], and LSeg [45]) that widely used in previous literature [14, 16, 15, 21]; 2) Instead of relying on tracking-based 2D methods for object counting in the scene [17], we exploit the multi-view global consistency of the 3D Gaussians to constrain the instance features. We adhere to the principle that Gaussian-rendered features from the same object should be close, while those from different objects should be distant. To achieve this, we employ binary SAM masks (instead of high-dimensional SAM features) without cross-view correlation to supervise the rendered instance feature maps using two types of losses: intra-mask smoothing loss and inter-mask contrastive loss.
@@
$\mathcal{L}_{c}=\frac{1}{m(m+1)}\sum_{i=1}^{m}\sum_{j=1,j\neq i}^{m}\frac{1}{\left\|\bar{\boldsymbol{M}}_{i}-\bar{\boldsymbol{M}}_{j}\right\|^{2}},$ (2)
@@ 3.2 Two-Level Codebook for Discretization
However, this solution still presents challenges: 1) Due to occlusions or distance, two objects may never share the same viewpoint and thus remain unoptimized by contrastive loss (i.e. Eq. (2)), failing to ensure their features are distinct. 2) In large scenarios, a $k$ value of 64 proves inadequate for distinguishing all objects, reducing the distinctiveness of instance features. However, Simply increasing $k$ offers limited benefits (will be demonstrated in experiments (Sec. 4.4)).
@@
$\left\{\begin{matrix}\left[\boldsymbol{F}\in\mathbb{R}^{n\times 6};\boldsymbol{X}\in\mathbb{R}^{n\times 3}\right]\mapsto\{\boldsymbol{C}_{\text{coarse}}\in\mathbb{R}^{k_{1}\times(6+3)},\boldsymbol{I}_{\text{coarse }}\in\{1,\ldots,k_{1}\}^{n}\}&\text{ coarse,}~{}k_{1}=64\\ \boldsymbol{F}\in\mathbb{R}^{n\times 6}\mapsto\{\boldsymbol{C}_{\text{fine }}\in\mathbb{R}^{(k_{1}\times k_{2})\times 6},\boldsymbol{I}_{\text{fine }}\in\{1,\ldots,k_{2}\}^{n}\}&\text{ fine,}~{}k_{2}=10\end{matrix}\right.$ (3)
@@
where $\boldsymbol{M_{p}}\in\mathbb{R}^{6\times H\times W}$ is the feature map rendered from the pseudo GT features, and $\boldsymbol{M_{c}}\in\mathbb{R}^{6\times H\times W}$ represents the feature map rendered from quantized features.
@@ 3.3 Instance-Level 2D-3D Association without Depth Test
4 Experiments
4.1 Open-Vocabulary Object Selection in 3D Space
Settings. 1) Task: Given an open-vocabulary text query, we extract its text feature using CLIP and calculate the cosine similarity between this feature and the language features of each Gaussian. Then, we select the highly relevant 3D points and render them into multi-view images using the 3DGS pipeline. 2) Baseline: We compared our method with LangSplat and LEGaussians. OpenGaussian associates each Gaussian with a 512-dimensional CLIP feature using the method described in Sec. 3.3. For LangSplat and LEGaussians, we followed their operation to reconstruct the 512-dimensional CLIP feature from the low-dimensional language feature of each Gaussian. 3) Dataset and Metrics: We conducted experiments on the Lerf-ovs dataset annotated by LangSplat. The average IoU and accuracy are calculated between the images rendered from the 3D Gaussian points selected by the text query and the GT object masks.
The visualization results are presented in Fig. 5. Given a query text, we can select the relevant Gaussian points and render them into multi-view images. However, the comparison methods make it difficult to identify the accurate target due to the ambiguous 3D point features. In the left two columns of Fig. 6, we show the results of feature visualization on the LERF dataset using PCA. Our features exhibit better discrimination.
Table 1: Performance of semantic segmentation on LeRF dataset compared to LangSplat and LEGaussians based on text query. Accurate is measured by mAcc@0.25.
@@ 4.1 Open-Vocabulary Object Selection in 3D Space
Methods figurines teatime ramen waldo_kitchen Mean figurines teatime ramen waldo_kitchen Mean
@@
4.2 Open-Vocabulary Point Cloud Understanding
@@ 4.4 Ablation Study
(1) The Necessity of the Two-Level Codebook. In case #1 of Tab 4, a single-layer codebook with $k=64$ was initially employed, resulting in a limited 28% mIoU. The primary limitation arose from the codebook’s capacity, which was insufficient to represent all objects in the scene. Increasing $k$ to $320$ in case #2 seemed like an intuitive solution, but it led to a significant decrease in performance. Visualizations highlighted that solely constraining instance features resulted in spatially distant points being grouped together within the same cluster. To address this, we introduced a two-level codebook approach. At the coarse level, we utilized both instance features and coordinates to ensure spatial proximity of 3D points within clusters. Then, at the fine level, we further discretized instance features. Notably, case #4 demonstrated substantial performance improvements with the two-level codebook. In contrast, case #3, which employed a two-level design without incorporating coordinates at the coarse level, underscored the positive impact of including coordinates.
(2) The Strategy for 2D-3D Feature Association. In our instance-level 2D-3D feature association, we assessed the IoU between 3D instance renderings and SAM masks, as well as the distance between instance features and pseudo-features. Through an ablation study presented in Tab 4, we found that each strategy can independently achieve comparable performance. Case #1 demonstrates the effectiveness of our codebook discretization, enabling accurate 3D instance acquisition to support the IoU-based association strategy. Case #2 highlights the discriminative power and global consistency of our instance features, showing that feature-based matching alone can effectively associate objects. Case #3 confirms that considering both strategies simultaneously yields the best performance.
@@
#3 ✓ ($k$=64) 30.27 46.44
#4 ✓ ($k$=64) 38.29 55.19
Table 4: Performance of semantic segmentation with different association strategies.
Case IoU Feat. dis. mIoU $\uparrow$ mAcc. $\uparrow$
#1 35.28 53.19
#2 34.01 51.35
#3 38.29 55.19
@@ 5 Conclusion
Appendix A Implementation Details
(1) Training Strategy. Consistent with LangSplat, we first pre-train the standard 3DGS for 30,000 steps. Subsequently, we freeze the Gaussian coordinates, scale, and opacity parameters, and train the instance features for 10,000 steps and the two-layer codebook for 20,000 steps. The 2D-3D feature association step is training-free. The extraction methods for SAM masks and CLIP features also align with LangSplat. While LangSplat extracts three layers of SAM masks (small, middle, and large), our implementation uses only one layer (large).
(2) Training time. We train each scene on a single 32G V100 GPU (with actual memory usage around 16 to 20G). For the LERF dataset, each scene takes around 200 images and trains for approximately 50 minutes. For the ScanNet dataset, each scene takes around 100-300 images, and trains for approximately 15 minutes. The 2D-3D feature association step is a one-time computation, and no further computation is needed during inference. The association process takes around 1 minute.
(3) ScanNet Dataset. We randomly selected 10 scenes from ScanNet for evaluation, specifically: scene0000_00, scene0062_00, scene0070_00, scene0097_00, scene0140_00, scene0200_00, scene0347_00, scene0400_00, scene0590_00, scene0645_00. The 19 categories (defined by ScanNet) used for text query are respectively: wall, floor, , cabinet, bed, chair, sofa, table, door, window, bookshelf, picture, counter, desk, curtain, refrigerator, shower curtain, toilet, sink, bathtub; 15 categories are without picture, refrigerator, showercurtain, bathtub; 10 categories are further without cabinet, counter, desk, curtain, sink.
Appendix B More Results
@@
Appendix C Splatting 3D Gaussian Points

### v2

@@ Abstract
This paper introduces OpenGaussian, a method based on 3D Gaussian Splatting (3DGS) capable of 3D point-level open vocabulary understanding. Our primary motivation stems from observing that existing 3DGS-based open vocabulary methods mainly focus on 2D pixel-level parsing. These methods struggle with 3D point-level tasks due to weak feature expressiveness and inaccurate 2D-3D feature associations. To ensure robust feature presentation and 3D point-level understanding, we first employ SAM masks without cross-frame associations to train instance features with 3D consistency. These features exhibit both intra-object consistency and inter-object distinction. Then, we propose a two-stage codebook to discretize these features from coarse to fine levels. At the coarse level, we consider the positional information of 3D points to achieve location-based clustering, which is then refined at the fine level. Finally, we introduce an instance-level 3D-2D feature association method that links 3D points to 2D masks, which are further associated with 2D CLIP features. Extensive experiments, including open vocabulary-based 3D object selection, 3D point cloud understanding, click-based 3D object selection, and ablation studies, demonstrate the effectiveness of our proposed method. The source code is available at our project page https://3d-aigc.github.io/OpenGaussian.
@@ 3 Method
3DGS [19] utilizes an explicit scene representation through 3D Gaussian points. Each Gaussian point encompasses various attributes such as position $\boldsymbol{\mu}$, rotation $\boldsymbol{R}$, scale $\boldsymbol{S}$, opacity $\sigma$, and spherical harmonics coefficients for representing direction-aware color $\boldsymbol{c}$. For a detailed description of the splatting process, please refer to Appendix A.3, where a set of 3D Gaussian points is projected onto 2D screen space and blended to generate pixels. Inspired from prior studies [33, 37, 52], we augment each 3D Gaussian point with a low-dimensional feature $\boldsymbol{f}\in\mathbb{R}^{6}$ to represent its instance attributes. However, our approach differs in two crucial aspects: 1) We do not require additional dimensional reduction, quantization, or distillation for pre-trained features (such as CLIP [34], SAM [21], DINO [7], and LSeg [22]) that widely used in previous literature [33, 52, 37, 8]; 2) Instead of relying on tracking-based 2D methods for object counting in the scene [48], we exploit the multi-view global consistency of the 3D Gaussians to constrain the instance features. We adhere to the principle that Gaussian-rendered features from the same object should be close, while those from different objects should be distant. To achieve this, we employ binary SAM masks (instead of high-dimensional SAM features) without cross-view correlation to supervise the rendered instance feature maps using two types of losses: intra-mask smoothing loss and inter-mask contrastive loss.
@@
$\mathcal{L}_{c}=\frac{1}{m(m-1)}\sum_{i=1}^{m}\sum_{j=1,j\neq i}^{m}\frac{1}{\left\|\bar{\boldsymbol{M}}_{i}-\bar{\boldsymbol{M}}_{j}\right\|^{2}},$ (2)
@@ 3.2 Two-Level Codebook for Discretization
However, this solution still presents challenges: 1) Due to occlusions or distance, two objects may never share the same viewpoint and thus remain unoptimized by contrastive loss (i.e. Eq. (2)), failing to ensure their features are distinct. 2) In large scenarios, a $k$ value of 64 proves inadequate for distinguishing all objects, reducing the distinctiveness of instance features. However, Simply increasing $k$ does not improve performance (will be demonstrated in experiments (Sec. 4.4)).
@@
$\left\{\begin{matrix}\left[\boldsymbol{F}\in\mathbb{R}^{n\times 6};\boldsymbol{X}\in\mathbb{R}^{n\times 3}\right]\mapsto\{\boldsymbol{C}_{\text{coarse}}\in\mathbb{R}^{k_{1}\times(6+3)},\boldsymbol{I}_{\text{coarse }}\in\{1,\ldots,k_{1}\}^{n}\}&\text{ coarse,}~{}k_{1}$=$64,32\\ \boldsymbol{F}\in\mathbb{R}^{n\times 6}\mapsto\{\boldsymbol{C}_{\text{fine }}\in\mathbb{R}^{(k_{1}\times k_{2})\times 6},\boldsymbol{I}_{\text{fine }}\in\{1,\ldots,k_{2}\}^{n}\}&\text{ fine,}~{}k_{2}$=$10,5\end{matrix}\right.$ (3)
@@
where $\boldsymbol{M_{p}}\in\mathbb{R}^{6\times H\times W}$ is the feature map rendered from the first stage pseudo features, and $\boldsymbol{M_{c}}\in\mathbb{R}^{6\times H\times W}$ represents the feature map rendered from quantized features.
@@ 3.3 Instance-Level 2D-3D Association without Depth Test
Table 1: Performance of object selection in 3D space from text query on LeRF dataset. Accurate is measured by mAcc@0.25. waldo_kitchen abbreviated as kitchen.
@@
Methods figurines teatime ramen kitchen Mean figurines teatime ramen kitchen Mean
@@
4 Experiments
4.1 Open-Vocabulary Object Selection in 3D Space
Settings. 1) Task: Given an open-vocabulary text query, we extract its text feature using CLIP and calculate the cosine similarity between this feature and the language features of each Gaussian. Then, we select the highly relevant 3D points and render them into multi-view images using the 3DGS pipeline. 2) Baseline: We compared our method with LangSplat and LEGaussians. OpenGaussian associates each Gaussian with a 512-dimensional CLIP feature using the method described in Sec. 3.3. For LangSplat and LEGaussians, we followed their operation to reconstruct the 512-dimensional CLIP feature from the low-dimensional language feature of each Gaussian. Note that our evaluations all follow a consistent setup: we use text to find matching 3D Gaussians, which are then rendered into multi-view images. Therefore, the metrics we report are inconsistent with the official metrics reported by comparison methods. 3) Dataset and Metrics: We conducted experiments on the Lerf-ovs dataset re-annotated by LangSplat. The average IoU and accuracy are calculated between the images rendered from the 3D Gaussian points selected by the text query and the GT object masks.
The visualization results are presented in Fig. 5. Given a query text, we can select the relevant Gaussian points and render them into multi-view images. However, the comparison methods make it difficult to identify the accurate target due to the ambiguous 3D point features. In the left two columns of Fig. 6, we show the results of feature visualization on the LERF dataset. Our features exhibit better discrimination.
@@ 4.4 Ablation Study
(1) Ablation of Intra-mask Smooth Loss and Inter-mask Contrastive Loss. To validate the effectiveness of the two losses proposed in Sec. 3.1, we conducted the ablation shown in Tab. 5. i) The inter-mask contrastive loss proves to be more crucial. Employing only this loss achieves respectable performance. Adding the intra-mask loss further enhances results, leading to a 3.05% improvement in mIoU and a 2.76% increase in mAcc. ii) The intra-mask smooth loss exhibits comparatively lower importance, which can be attributed to the inherent characteristics of 3DGS, where a single Gaussian point represents multiple pixels. Consequently, features of neighbouring pixels tend to be similar, indicating that 3DGS naturally induces a smoothing effect on adjacent pixels. This intrinsic smoothing mechanism partially mitigates the contribution of the intra-mask smooth loss.
(2) The Necessity of the Two-Level Codebook. i) In case #1 of Tab 5, a single-layer codebook with $k=64$ was initially employed, resulting in a limited 28% mIoU. The primary limitation arose from the codebook’s capacity, which was insufficient to represent all objects in the scene. Increasing $k$ to $320$ in case #2 seemed like an intuitive solution, but it led to a significant decrease in performance. Visualizations highlighted that solely constraining instance features resulted in spatially distant points being grouped together within the same cluster. ii) To address this, we introduced a two-level codebook approach. At the coarse level, we utilized both instance features and coordinates to ensure spatial proximity of 3D points within clusters. Then, at the fine level, we further discretized instance features. Notably, case #5 demonstrated substantial performance improvements with the two-level codebook. In contrast, case #6, which employed a two-level design without incorporating coordinates at the coarse level, underscored the positive impact of including coordinates. iii) To illustrate the importance of considering both coordinates and the two-level codebook, we conducted experiments with the case #3 and case #4 configurations, in which only position information was considered without using the two-level codebook.
(3) The Strategy for 2D-3D Feature Association. In our instance-level 2D-3D feature association, we assessed the IoU between 3D instance renderings and SAM masks, as well as the distance between instance features and pseudo-features. Through an ablation study presented in Tab 5, we found that each strategy can independently achieve comparable performance. Case #1 demonstrates the effectiveness of our codebook discretization, enabling accurate 3D instance acquisition to support the IoU-based association strategy. Case #2 highlights the discriminative power and global consistency of our instance features, showing that feature-based matching alone can effectively associate objects. Case #3 confirms that considering both strategies simultaneously yields the best performance. All the ablations are evaluated on the semantic segmentation task of the 10 categories on ScanNet.
Table 3: Inter/Intra loss ablation.
Case Inter Intra mIoU $\uparrow$ mAcc. $\uparrow$
#1 35.24 52.43
#2 25.89 42.76
#3 38.29 55.19
Table 4: Association strategy ablation.
Case IoU Feat. dis. mIoU $\uparrow$ mAcc. $\uparrow$
#1 35.28 53.19
#2 34.01 51.35
#3 38.29 55.19
@@
#3 ✓ ($k$=64) 32.04 49.82
#4 ✓ ($k$=320) 15.20 24.91
#5 ✓ ($k$=64) 30.27 46.44
#6 ✓ ($k$=64) 38.29 55.19
@@ 5 Conclusion
Limitations: (1) The geometric properties of the Gaussian (position, opacity, scale) are fixed. This may lead to inconsistencies between geometric representation and semantic content. We will consider joint optimization of instance features and geometric properties in future work. (2) The values of $k$ for the two-level codebooks are determined empirically. It is necessary to study scenario-specific adaptive values to optimize performance across diverse contexts. (3) We focus on 3D point-level understanding without considering the regression of object sizes to perform open-vocabulary 3D detection tasks [27, 6, 5, 46]. (4) Currently, we have not considered dynamic factors, which are common challenges in real-world applications. Integrating the proposed method with 4DGS [43, 13] would be meaningful.
Appendix A Appendix
A.1 Implementation Details
(1) Training Strategy. Consistent with LangSplat, we first pre-train the standard 3DGS for 30,000 steps. Subsequently, we freeze the Gaussian coordinates, scale, and opacity parameters, and train the instance features for 10,000 steps (ScanNet is 20,000 steps) and the two-layer codebook for 30,000 steps (ScanNet is 40,000 steps). The 2D-3D feature association step is training-free. The extraction methods for SAM masks and CLIP features also align with LangSplat. While LangSplat extracts three layers of SAM masks (small, middle, and large), our implementation uses only one layer (large).
(2) Training Time. We train each scene on a single 32G V100 GPU (with actual memory usage around 16 to 20G). For the LERF dataset, each scene takes around 200 images and trains for approximately 50 minutes. For the ScanNet dataset, each scene takes around 100-300 images (Sample every 20 frames from the original data and downsample by 2), and trains for approximately 15 minutes. The 2D-3D feature association step is a one-time computation, and no further computation is needed during inference. The association process takes around 1 minute.
(3) ScanNet Dataset Evaluation. We randomly selected 10 scenes from ScanNet for evaluation, specifically: scene0000_00, scene0062_00, scene0070_00, scene0097_00, scene0140_00, scene0200_00, scene0347_00, scene0400_00, scene0590_00, scene0645_00.
The 19 categories (defined by ScanNet) used for text query are respectively: wall, floor, cabinet, bed, chair, sofa, table, door, window, bookshelf, picture, counter, desk, curtain, refrigerator, shower curtain, toilet, sink, bathtub;
15 categories are without picture, refrigerator, showercurtain, bathtub;
10 categories are further without cabinet, counter, desk, curtain, sink.
(4) Hyperparameters. 1) The values of $k$ in the two-level codebook. In the ScanNet dataset, $k1=64,k2=5$ are used uniformly. In the LeRF dataset, for the teatime scene, $k1=32,k2=10$; for the other scenes, $k1=64,k2=10$. 2) The weights of the coordinates in the coarse-level codebook. In the ScanNet dataset, the weight is $1.0$. In the LeRF dataset, the weight for the teatime scene is $0.1$, while for the other scenes, the weight is $0.5$. 3) The weight of the intra-mask smoothing loss. In the ramen scene of LeRF, the weight is $0.01$; for the other scenes and ScanNet, the weight is $0.1$.
A.2 More Results
A.2.1 Scene editing
@@ A.2.2 Instance feature visualization
A.2.3 Qualitative results of outdoor and real-world scenarios
Fig. 13 shows qualitative results of 3D instance features for 6 sequences in the Waymo outdoor dataset, demonstrating the capability to discretize large cases.
Fig. 14 presents the results of rendering 3D features into 2D feature maps, showcasing the ability to learn instance features with 3D consistency from coarse SAM supervision.
Fig. 15 illustrates the effectiveness of the two-stage codebook in outdoor scenes.
Furthermore, we conducted validations in the real-world scene. As depicted in Fig. 11, using an office scene captured by a mobile phone, the visualization of 3D points demonstrates that OpenGaussian achieved significant object discrimination in the real world.
A.2.4 Text-to-3D Gaussian retrieval
Fig. 9 shows a demo of retrieving relevant Gaussians via text query, which is achieved by computing the cosine similarity between text features and the language features associated with 3D Gaussians (Sec. 3.3).
@@ A.2.4 Text-to-3D Gaussian retrieval
Figure 9: A demo of text-to-3D Gaussian retrieval on ScanNet. Top: scene0000_00; bottom: scene0645_00.
@@
Figure 11: Visualization of 3D point features in the real-world scene captured by mobile phone.
@@
Figure 13: Feature visualization of 3D points on the large-scale outdoor dataset Waymo. (a)-(f) are 6 different scenes selected from the Waymo dataset. Left: RGB image; Right: 3D point features.
Figure 14: Results of rendering 3D point features onto images in the Waymo dataset. We trained 3D point features with multi-view consistency using SAM masks that without inter-frame associations.
Figure 15: Validation of two-level codebook in outdoor scenes. Achieved better discretization with the fine-level codebook.
A.3 Splatting 3D Gaussian Points